# Learning Robust Intervention Representations with Delta Embeddings

### v1

@@ Abstract
Causal representation learning has attracted significant research interest during the past few years, as a means for improving model generalization and robustness. Causal representations of interventional image pairs, have the property that only variables corresponding to scene elements affected by the intervention / action are changed between the start state and the end state. While most work in this area has focused on identifying and representing the variables of the scene under a causal model, fewer efforts have focused on representations of the interventions themselves. In this work, we show that an effective strategy for improving out of distribution (OOD) robustness is to focus on the representation of interventions in the latent space. Specifically, we propose that an intervention can be represented by a Causal Delta Embedding that is invariant to the visual scene and sparse in terms of the causal variables it affects. Leveraging this insight, we propose a framework that is capable of learning causal representations from image pairs, without any additional supervision. Experiments in the Causal Triplet challenge demonstrate that Causal Delta Embeddings are highly effective in OOD settings, significantly exceeding baseline performance in both synthetic and real-world benchmarks.
@@ Introduction
Understanding and reasoning about how the world changes in response to actions and external interventions is fundamental for artificial intelligence (AI) agents, especially those operating in dynamic environments. While deep learning models excel at learning complex patterns from data, they often struggle to generalize to new situations where the underlying data distribution changes, which is a critical limitation for real world deployment (Hendrycks et al. 2021; Geirhos et al. 2020). Recovering the true mechanisms that govern how data is generated and transformed is important for intelligent agents that can reason causally, as opposed to those that rely on statistical correlations (Pearl 2009).
This fundamental problem falls under the category of Causal Representation Learning (CRL) (Schölkopf et al. 2021), which focuses on disentangling the underlying causal variables of a system (Khemakhem et al. 2020). However, despite its importance in practical applications such as robotics or healthcare (Gupta et al. 2024; Hellström 2021; Sanchez et al. 2022), the specific challenge of learning robust, generalizable representations of the causal variables, as well as the interventions themselves remains an open problem.
@@
A core challenge in building such robust causal AI systems lies in faithfully capturing the underlying data generation process. This is guided by two fundamental principles within CRL. First, the Independent Causal Mechanisms (ICM) posits that the distribution’s generative process can be decomposed into autonomous and independent modules, each representing a distinct causal mechanism (Peters, Janzing, and Schölkopf 2017; Schölkopf et al. 2021). Second, the Sparse Mechanism Shift (SMS) assumption suggests that an intervention typically affects only a small, localized subset of these causal mechanisms (Schölkopf et al. 2021). Most existing methods focus on identifying these disentangled factors from observations (Higgins et al. 2017; Khemakhem et al. 2020; Ahuja, Hartford, and Bengio 2022). Fewer methods, however, have focused on learning generalizable representations of the interventions, which can be equally important in predicting the outcome of interventions, especially when faced with novel situations.
@@
In this paper, we introduce Causal Delta Embedding (CDE), a novel framework for learning robust representations of interventions from image pairs. Our core insight is that the intervention can be effectively isolated and represented as a delta embedding, the vector difference between the latent representations of pre- and post- intervention states if it satisfies the properties of (a) independence to causally irrelevant elements of the scene, in accordance to the ICM assumption (b) sparsity, in accordance to the SMS assumption and (c) object invariance, i.e., that the representation remains the same across objects. Using these properties as a guide, a learning strategy is proposed for learning CDEs from interventional image pairs.
@@
We evaluate CDE on the Causal Triplet challenge (Liu et al. 2023), which encompasses 3 increasingly complex settings: single-object synthetic data, multi-object synthetic data and real world scenes from EpicKitchens (Damen et al. 2022). Our experiments demonstrate that CDE establishes a new state of the art in OOD generalization for this challenge. Beyond quantitative performance, our analysis reveals that CDE learns a semantically structured intervention space, autonomously discovering anti-parallel relationships between opposing actions (e.g., open vs. close) without any explicit supervision.
@@
  • We introduce Causal Delta Embedding (CDE), a novel theoretical framework that explicitly defines the desiderata and learning principles for robustly representing interventions in a disentangled latent space.
@@
Part of the research on CRL focuses on identifying latent causal variables from high dimensional observations (Khemakhem et al. 2020; Ahuja, Hartford, and Bengio 2022). These methods established identifiablity conditions for nonlinear ICA and demonstrated causal factor recovery under specific assumptions (Wendong et al. 2023; Monti, Zhang, and Hyvärinen 2020). Another line of work focuses on causal disentanglement (Yang et al. 2021; Shen et al. 2020; Brehmer et al. 2022; Locatello et al. 2020a). These approaches often extend the Variational Autoencoder (VAE) framework (Kingma, Welling et al. 2013; Higgins et al. 2017). Object-centric learning methods have also been proposed in order to disentangle the visual scene into manipulable objects (Locatello et al. 2020b; Seitzer et al. 2022). More recent work has leveraged interventional data to improve causal disentanglement (Brehmer et al. 2022; Lippe et al. 2022; Squires et al. 2023; Lippe et al. 2023; Ahuja, Hartford, and Bengio 2022), showing that interventions provided crucial inductive biases for learning causal representations (Ahuja et al. 2023). While previous methods focus on identifying causal variables, our work instead models the interventional mechanisms, by learning generalizable embeddings that represent interventions in a way that remains invariant across different contexts.
@@
Traditional action recognition methods rely on spatiotemporal patterns and achieve strong performance under IID conditions (Carreira and Zisserman 2017; Feichtenhofer et al. 2019; Arnab et al. 2021). However, these correlation-based approaches struggle with distribution shifts (Geirhos et al. 2020) and often are associated with spurious correlations (Wang and Jordan 2024). Recent work has explored domain adaptation (Chen et al. 2019; Munro and Damen 2020) and causal approaches (Magliacane et al. 2018; Wang et al. 2023) for robust action understanding. Another category of methods uses large Vision Language Action (VLA) models (Kim et al. 2024; Zitkovich et al. 2023; Ma et al. 2024) to enable agents to perform actions in challenging environments. These models typically depend on large-scale pre-training on diverse data, yet generalization to unseen tasks remains an open challenge (Sapkota et al. 2025). Unlike these approaches, our method learns causal representations of interventions, in the sense that they satisfy properties resulting from the CRL assumptions. These representations are shown to generalize to novel object-action combinations without the need for finetuning.
@@
Contrastive learning has proven effective for learning meaningful representations by contrasting similar and dissimilar examples (Chen et al. 2020; Khosla et al. 2020). However, existing methods contrast individual samples, rather than relationships between samples. The principle of sparse mechanism shifts (Schölkopf et al. 2021; Peters, Janzing, and Schölkopf 2017), suggests that interventions affect only a small subset of the causal mechanisms. Sparsity in causal representations has been explored in various works (Pfister and Peters 2022; Xu et al. 2024) and has shown to improve disentanglement. However, combining sparsity with adversarial training (Liu et al. 2023) often leads to poor OOD performance, since other confounders might still be present in the scene, motivating our approach for stricter assumptions.
Problem Formulation
@@
We formalize this challenge within the framework of the Structural Causal Model presented in (Liu et al. 2023) (Figure 2). Let us consider a set of causal variables $Z\in\mathcal{Z}\subset\mathbb{R}^{l}$, representing the state of the underlying data generating mechanisms. These variables have dependencies that are defined through a set of structural equations:
$Z_{i}:=f_{i}(\text{pa}(Z_{i}),\epsilon_{i}),\quad i=1,\ldots,l$
where $\text{pa}(Z_{i})$ denotes the set of causal parents of variable $Z_{i}$, and the $\epsilon_{i}$ are mutually independent stochastic noise terms representing unmodeled factors. The high-dimensional visual observation $x\in\mathcal{X}\subset\mathbb{R}^{d}$ is rendered from these latent variables via a complex, non-invertible generative function $g:\mathcal{Z}\to\mathcal{X}$, such that $x=g(Z)$.
We assume the latent space $Z$ can be partitioned into scene-level variables $Z_{s}$ (e.g., illumination, camera pose) and a set of object-level variables $Z_{o}=\{Z_{n,k}\}_{n=1,k=1}^{N,K}$, corresponding to the $k$-th property of the $n$-th object. An agent performs an action $a\in\mathcal{A}$, which performs an intervention on the system. This intervention transforms the pre-intervention state $Z$ into a post-intervention state $\tilde{Z}$.
Unobserved confounders ($c$) create spurious correlations and a training-testing distribution mismatch $P_{\text{train}}(Z,a)\neq P_{\text{test}}(Z,a)$. Following the Independent Causal Mechanisms principle (Schölkopf et al. 2012; Peters, Janzing, and Schölkopf 2017), we assume the true causal mechanism $P(\tilde{Z}_{a}|Z_{a},a)$ is invariant to this shift. Therefore, a robust model must learn this invariant mechanism instead of non-stationary correlations.
@@
Figure 2: Causal Graph for a pair of observations (x,x~)($\tilde{x}$) before and after an action $a$, proposed in (Liu et al. 2023). The data generating process is described by a set of latent factors, including global scene level factors $z_{s}$ and local object level factors $z_{n}^{k}$, which are dependent due to confounders $c$. The action is assumed to influence only a few object level causal factors $z_{a}$ in the scene and the effect of that influence is captured by $\tilde{z}_{a}$. The red dashed line indicates the structural equation assumed by our CDE approach.
@@
Causal Delta Embeddings
@@
If the encoder is faithful to the data generating process illustrated by the model of Figure 2 and assuming identical noise across observations then for the Delta Embedding we have
$\delta_{a}$ $=\begin{bmatrix}z_{s}&z_{1}&\cdots&\tilde{z}_{a}&\cdots z_{N}\end{bmatrix}^{T}$-
$\begin{bmatrix}z_{s}&z_{1}&\cdots&z_{a}&\cdots z_{N}\end{bmatrix}^{T}$
$=\begin{bmatrix}0&\cdots&\tilde{z}_{a}-z_{a}&\cdots&0\end{bmatrix}^{T}$ (1)
where $z_{a}$ is the dimension (or subset of dimensions) of object $n$ that is affected by action $a$. From Eq. (1) we observe the following properties of $\delta_{a}$.
@@
    Independence. Under the model of Figure 2, an action’s representation is independent of the causally irrelevant elements of the scene, i.e. scene properties and objects not affected by $a$.
@@
    Sparsity. If the assumption of Sparse Mechanism Shifts (Schölkopf et al. 2021; Liu et al. 2023) holds, then the action $a$ will affect only a few underlying causal factors of the system, and the representation of the change, $\delta_{a}$, will be sparse.
@@
$\tilde{z}_{a}=f(z_{a},\epsilon)=z_{a}+\delta_{a}+\mathcal{N}(0,\sigma\mathbb{I})$ (3)
where $\sigma$ is small. The following section uses this definition to develop a strategy for learning Causal Delta Embeddings.
Approach
The Global Causal Delta Embedding Model
We first introduce a global model, i.e., a model that learns a single causal representation from the entire image.
Model Architecture
The model consists of three main components, as illustrated in Figure 3 (A).
@@
  1. 1.
    The Encoder ($\phi$): The encoder is responsible for mapping an input image $x$ into the $\mathcal{Z}$. It is composed of two sub-modules:
    • A Pre-trained Vision Backbone: We use a powerful Vision Transformer (ViT) (Dosovitskiy et al. 2020), specifically one pre-trained with the DINO self-supervision algorithm (Caron et al. 2021). The backbone processes the input image and outputs a high-dimensional feature vector. We use the output corresponding to the ‘[CLS]’ token as the global image representation.
    • A Causal Projector: The feature vector from the backbone is then passed through a small multi-layer perceptron (MLP). This projector’s role is to transform the general-purpose features into an $l$-dimensional representation satisfying the Causal Delta Embedding properties.
  2. 2.
@@
    $\delta=\tilde{z}-z$
  3. 3.
    The Action Classifier ($h$): The resulting delta vector $\delta$ is the sole input to a final classification head. This head is another small MLP that takes the $l$-dimensional delta and outputs logits for the different action classes in $\mathcal{A}$.
Implementation of the Learning Objective
To train the model and learn the Causal Delta Embeddings satisfying the properties outlined in the previous section, we combine three loss functions.
  1. 1.
    Cross-Entropy Loss: The primary objective is to ensure the delta embedding is useful for the downstream task. We use a standard Cross-Entropy loss, $\mathcal{L}_{\text{CE}}$ between the predicted action logits $h(\delta_{i})$ and the one-hot ground-truth action label $a_{i}$.
  2. 2.
    Supervised Contrastive Loss: To enforce that embeddings for the same action are clustered together (Property 3, Invariance), we employ the Supervised Contrastive Loss, $\mathcal{L}_{\text{contrast}}$ (Khosla et al. 2020). For a batch of $B$ delta embeddings, the loss for each embedding $\delta_{i}$ (the “anchor”) encourages it to be closer to other embeddings of the same class (“positives”) than to all other embeddings in the batch.

### v2

@@ Abstract
Causal representation learning has attracted significant research interest during the past few years, as a means for improving model generalization and robustness. Causal representations of interventional image pairs (also called “actionable counterfactuals” in the literature), have the property that only variables corresponding to scene elements affected by the intervention / action are changed between the start state and the end state. While most work in this area has focused on identifying and representing the variables of the scene under a causal model, fewer efforts have focused on representations of the interventions themselves. In this work, we show that an effective strategy for improving out of distribution (OOD) robustness is to focus on the representation of actionable counterfactuals in the latent space. Specifically, we propose that an intervention can be represented by a Causal Delta Embedding that is invariant to the visual scene and sparse in terms of the causal variables it affects. Leveraging this insight, we propose a method for learning causal representations from image pairs, without any additional supervision. Experiments in the Causal Triplet challenge demonstrate that Causal Delta Embeddings are highly effective in OOD settings, significantly exceeding baseline performance in both synthetic and real-world benchmarks.
@@ 1 Introduction
Understanding how the world changes in response to actions and external interventions is fundamental for artificial intelligence agents, especially those operating in dynamic environments. Although deep learning models are highly successful at capturing complex patterns from data, they often fail to generalize to new situations where the underlying data distribution changes, which is a critical limitation for real world deployment Hendrycks et al. (2021); Geirhos et al. (2020). To overcome this, agents must recover the underlying mechanisms that generate and transform data, enabling causal reasoning and robust generalization (Pearl, 2009).
@@
This fundamental problem falls within the scope of Causal Representation Learning (CRL) (Schölkopf et al., 2021), which seeks to disentangle the causal variables of a system (Khemakhem et al., 2020a). Despite its importance in practical applications such as robotics or healthcare (Gupta et al., 2024; Hellström, 2021; Sanchez et al., 2022; Tejada-Lapuerta et al., 2023), the challenge of learning disentangled and generalizable representations of the causal variables remains unresolved. Addressing this challenge requires accurate modelling of the underlying data generation process, a task which is guided by two fundamental assumptions within CRL. First, the Independent Causal Mechanisms (ICM) assumption, which posits that the distribution’s generative process can be decomposed into autonomous and independent modules, each representing a distinct causal mechanism (Peters et al., 2017; Schölkopf et al., 2021). Second, the Sparse Mechanism Shift (SMS) assumption, which suggests that an intervention typically affects only a small, localized subset of these causal mechanisms (Schölkopf et al., 2021). While foundational methods focused on identifying these mechanisms from observational data or weak supervision (Brehmer et al., 2022; Lippe et al., 2022; Buchholz et al., 2023), recent advances have successfully leveraged interventional data to establish stronger identifiability guarantees, utilizing score-based methods (Kulkarni et al., 2025), unknown multi-node interventions (Varıcı et al., 2024), or invariance principles (Yao et al., 2024). However, the primary focus of these approaches remains on recovering the latent state variables. Fewer methods have focused on learning generalizable representations of actions (interventions), which can be equally important in predicting the outcome of interventions, especially when faced with novel situations.
@@
In this paper, we introduce Causal Delta Embedding (CDE), a novel framework for learning robust representations of interventions from image pairs. Using CDEs the intervention can be effectively isolated and represented as the vector difference between the latent representations of pre- and post- intervention states if it satisfies the properties of (a) independence to causally irrelevant elements of the scene, in accordance to the ICM assumption (b) sparsity, in accordance to the SMS assumption and (c) object invariance, i.e., that the representation remains the same across objects. Using these properties as a guide, a learning strategy is proposed for learning CDEs from interventional image pairs.
@@
We evaluate CDE on the Causal Triplet challenge (Liu et al., 2023), which encompasses 3 increasingly complex settings: single-object synthetic data, multi-object synthetic data and real world scenes from Epic Kitchens (Damen et al., 2022). Our experiments demonstrate that CDE establishes a new state of the art in OOD generalization for this challenge. Beyond quantitative performance, our analysis reveals that CDE learns semantically meaningful representations in the intervention space, autonomously discovering anti-parallel relationships between opposing actions (e.g., open vs. close) without any explicit supervision.
@@
  • We introduce Causal Delta Embedding (CDE), a novel approach for learning generalizable representations of interventions in a disentangled latent space.
@@ 2 Related Work
Research on CRL spans multiple directions. One line of work focuses on identifying latent causal variables from high-dimensional observations (Khemakhem et al., 2020a). These methods established identifiability conditions for nonlinear ICA and demonstrated causal factor recovery under specific assumptions (Wendong et al., 2023; Monti et al., 2020; Khemakhem et al., 2020b). Recent theoretical advances leverage score-based methods to achieve identifiability even with unknown multi-node interventions (Varıcı et al., 2024; Varici et al., 2023), while practical applications demonstrate use cases in robotics (Kulkarni et al., 2025) and single-cell genomics (Lopez et al., 2023). Another body of work focuses on causal disentanglement (Yang et al., 2021; Shen et al., 2020; Locatello et al., 2020a), often extending the VAE framework (Kingma et al., 2013; Higgins et al., 2017) and leveraging interventional data (Brehmer et al., 2022; Lippe et al., 2022; Squires et al., 2023; Lippe et al., 2023; Ahuja et al., 2022; 2023; Buchholz et al., 2023) while other methods focus on object-centric representations in order to disentangle visual scenes into manipulable objects (Locatello et al., 2020b; Seitzer et al., 2022). The invariance principle has emerged as a unifying framework (Yao et al., 2024), showing that many CRL methods exploit distributional symmetries created by interventions rather than requiring explicit causal semantics. While most of these methods focus on recovering causal variables and their relationships, our work takes a fundamentally different approach: we learn embeddings that represent the interventions themselves, capturing how mechanisms change rather than identifying which variables exist.
@@
Traditional action recognition methods rely on spatiotemporal patterns and achieve strong performance under IID conditions (Carreira and Zisserman, 2017; Feichtenhofer et al., 2019; Arnab et al., 2021). However, these correlation-based approaches struggle with distribution shifts (Geirhos et al., 2020) and often rely on spurious correlations (Wang and Jordan, 2024). Recent work has explored domain adaptation (Chen et al., 2019; Munro and Damen, 2020) and causal approaches (Magliacane et al., 2018; Wang et al., 2023) for robust action understanding. Another category of methods uses large Vision Language Action (VLA) models (Kim et al., 2024; Zitkovich et al., 2023; Ma et al., 2024) to enable agents to perform actions in challenging environments. These models typically depend on large-scale pre-training on diverse data, yet generalization to unseen tasks remains an open challenge (Sapkota et al., 2025). Unlike these approaches, our method learns causal representations of interventions that satisfy properties resulting from CRL assumptions, generalizing to novel object-action combinations without fine-tuning.
@@
Contrastive learning effectively learns meaningful representations by contrasting similar and dissimilar examples (Chen et al., 2020; Khosla et al., 2020; He et al., 2020; Grill et al., 2020). Recent theoretical work establishes deep connections between contrastive learning and causal structure discovery. Multi-view contrastive methods handle partial observability by learning shared representations across modalities (Yao et al., 2023; Federici et al., 2020; Tian et al., 2020). When data augmentations correspond to causal interventions, contrastive learning can provably disentangle causal factors (Von Kügelgen et al., 2021; Zimmermann et al., 2021). However, these methods contrast individual samples or augmented views, whereas our approach contrasts relationships between pre- and post-intervention pairs aiming at learning embeddings of the transformations themselves rather than sample level representations. The sparse mechanism shift (SMS) principle (Schölkopf et al., 2021; Peters et al., 2017) suggests interventions affect only small subsets of causal mechanisms. This sparsity assumption enables identifiability in various settings, such as mechanism sparsity regularization for nonlinear ICA (Lachapelle et al., 2022; 2024) or instance dependent partial observability (Xu et al., 2024). Recent work demonstrates that combining sparsity with appropriate regularization improves both identifiability (Pfister and Peters, 2022; Buchholz et al., 2023) and generalization (Lopez et al., 2023; Layne et al., 2024). While adversarial training with sparsity constraints has been explored (Liu et al., 2023), it often yields poor OOD performance when confounders remain in the scene, motivating our stricter causal assumptions and explicit intervention embedding approach.
3 Problem Formulation
@@
We formalize this challenge within the framework of the Structural Causal Model presented by Liu et al. (2023) (Figure 2). Let us consider a set of causal variables $z\in\mathcal{Z}\subset\mathbb{R}^{l}$, representing the state of the underlying data generating mechanisms. These variables have dependencies that are defined through a set of structural equations. Assuming an additive noise model, these equations are of the form
$z_{i}:=f_{i}(\text{pa}(z_{i}))+\epsilon_{i},\quad i=1,\ldots,l$
where $\text{pa}(z_{i})$ denotes the set of causal parents of variable $z_{i}$, and the $\epsilon_{i}$ are mutually independent stochastic noise terms representing unmodeled exogenous factors. The high-dimensional visual observation $x\in\mathcal{X}\subset\mathbb{R}^{d}$ is rendered from these latent variables via a complex generative function $g:\mathcal{Z}\to\mathcal{X}$, such that $x=g(z)$.
We assume the latent variable $z$ can be partitioned into scene-level variables $z_{s}$ (e.g., illumination, camera pose) and a set of object-level variables $z_{o}=\{z_{n,k}\}_{n=1,k=1}^{N,K}$, corresponding to the $k$-th property of the $n$-th object. An agent performs an action $a\in\mathcal{A}$, which applies an intervention on the system. This intervention transforms the pre-intervention state $z$ into a post-intervention state $\tilde{z}$. Unobserved confounders ($c$) create spurious correlations and a training-testing distribution mismatch $P_{\text{train}}(Z=z,a)\neq P_{\text{test}}(Z=z,a)$. Following the Independent Causal Mechanisms principle (Schölkopf et al., 2012; Peters et al., 2017), we assume the true causal mechanism $P(\tilde{Z}_{a}=\tilde{z}_{a}|Z_{a}=z_{a},a)$ is invariant to this shift. Therefore, a robust model must learn this invariant mechanism instead of non-stationary correlations.
@@
Figure 2: Causal Graph for a pair of observations (x,x~)($\tilde{x}$) before and after an action $a$, proposed by Liu et al. (2023). The data generating process is described by a set of latent factors, including global scene level factors $z_{s}$ and local object level factors $z_{n}^{k}$, which are dependent due to confounders $c$. The action is assumed to influence only a few object level causal factors $z_{a}$ in the scene and the effect of that influence is captured by $\tilde{z}_{a}$. The red dashed line indicates the structural equation assumed by our CDE approach.
@@
Notice also that the exogenous noise $\epsilon$ may differ across states $z$ and $\tilde{z}$, especially in real-world data, where there is no control over the environment and data acquisition conditions. For this reason, they are referred to as “actionable counterfactuals” in Liu et al. (2023) (as opposed to perfect counterfactuals).
@@
4 Causal Delta Embeddings
@@
Assuming identical noise across observations (i.e., perfect counterfactuals, the actionable counterfactual case will be discussed in the following section) and that $z_{a}$ in the data generating process of Figure 2 is identifiable by the encoder, for the Delta Embedding we have
$\delta_{a}$ $=\begin{bmatrix}0&\cdots&\tilde{z}_{a}-z_{a}&\cdots&0\end{bmatrix}^{T}$ (1)
where $z_{a}$ is the dimension (or subset of dimensions) of object $n$ that is affected by action $a$ in accordance with the notation used in the model of Figure 2. From equation 1 we observe the following properties of $\delta_{a}$.
@@
    Independence. Under the model of Figure 2, an action’s representation is independent of scene properties and objects not affected by $a$, i.e., it is not influenced or informed by them.
@@
    Sparsity. If the assumption of Sparse Mechanism Shifts Schölkopf et al. (2021); Liu et al. (2023) holds, then the action $a$ will affect only a few underlying causal factors of the system, and the representation of the change, $\delta_{a}$, will be sparse.
@@
$\tilde{z}_{a}=f(z_{a},a,\epsilon)=z_{a}+\delta_{a}+\epsilon$ (3)
where $\epsilon$ is the value of zero mean, independent exogenous noise. When the only change before and after the intervention is the effect of action $a$ (perfect counterfactual), $\epsilon=0$. The following section uses this definition to develop a strategy for learning Causal Delta Embeddings.
5 Approach
5.1 The Global Causal Delta Embedding Model
5.1.1 Model Architecture
We first introduce a global model, i.e., a model that learns a single causal representation from the entire image. The model consists of three main components, as illustrated in Figure 3 (A).
@@ 5.1.1 Model Architecture
The Encoder ($\phi$):
The encoder is responsible for mapping an input image $x$ into the $\mathcal{Z}$. It is composed of two sub-modules. (i) A Pre-trained Vision Backbone: We use a powerful Vision Transformer (ViT) (Dosovitskiy et al., 2020), specifically one pre-trained with the DINO self-supervision algorithm (Caron et al., 2021). The backbone processes the input image and outputs a high-dimensional feature vector. We use the output corresponding to the ‘[CLS]’ token as the global image representation. (ii) A Causal Projector: The feature vector from the backbone is then passed through a small multi-layer perceptron (MLP). This projector’s role is to transform the general-purpose features into an $l$-dimensional representation satisfying the Causal Delta Embedding properties.
Delta Computation and the action classifier $h$:
@@
5.1.2 Implementation of the Learning Objective
To learn CDEs that satisfy the properties outlined in the Section 4, we combine three loss functions. (i) Cross-Entropy Loss: The primary objective is to ensure the delta embedding is useful for the downstream task. We use a standard Cross-Entropy loss, $\mathcal{L}_{\text{CE}}$ between the predicted action logits $h(\delta_{i})$ and the one-hot ground-truth action label $a_{i}$. (ii) Supervised Contrastive Loss: To learn embeddings that are clustered together for the same action (Property 3, Invariance), we use the Supervised Contrastive Loss, $\mathcal{L}_{\text{contrast}}$ (Khosla et al., 2020). For a batch of $B$ delta embeddings, the loss for each embedding $\delta_{i}$ (the “anchor”) encourages it to be closer to other embeddings of the same class (“positives”) than to all other embeddings in the batch.